# Understanding Codebase like a Professional! Human–AI Collaboration for Code Comprehension

### v2

@@ 6. Stage 3: Prototype Design
We designed a prototype, CodeMap, to operationalize the four design opportunities.
@@ 6.1. Usage Scenario
We illustrate CodeMap with a concrete scenario in Figure 3 and a demonstration video on our website (Authors, [n. d.]). Suppose a novice developer, Bob, needs to understand a new codebase. While reading the source code in VSCode, he uploads the repository to CodeMap. After a brief waiting period, CodeMap generates a high-level business component map of the project. Next, Bob reads the global understanding project overview and follows step-by-step guidance to locate the codebase’s entry point and main modules (Figure 4). Meanwhile, he switches back to VSCode to inspect the original code and verify whether the relationships shown in the map.
@@
Figure 3. Interface for an example business component map. To support codebase reading, CodeMap allows users to start with 1) Global Understanding through the Global Business Component Map, which provides natural language–enhanced guidance on key modules, components, files, functions and their relationships. Then, users can 2) zoom into a specific module (e.g., “Security Module”) to view textual instructions in the right pane that describe the module’s details and its relationships with other modules.
Figure 4. An illustration of how the user uses the system.
@@ 6.2. Key Features
  • Global map: business component map: extracting high-level structural information, including business modules, components, and links to show their relationships; function call map: mainly visualizing classes as nodes and showing their inheritance relationships.
@@
  • Natural language enhancement for readability: presenting the links between nodes in the map and natural language to describe the relationships between modules. For example, “Component A provides or manages data for Component B”. Moreover, providing step-by-step guidance for global map for logically orienting users through project’s modules.
In this way, users no longer need to gather information manually and can focus their efforts on understanding the overall structure and organization of the codebase.
Key Feature 2: Seamless switching between layers by interaction.
@@ 6.4. Accuracy Reliability Checking
Iterative Refinement. We implemented an iterative refinement process—commonly used for LLM-generated outputs (Gu et al., 2024; Liu et al., 2024)—for graph generation; each iteration builds upon the previous iteration’s output. The iteration prompts are available on our website.
Accuracy Evaluation. There is no definitive ground truth for the business component map, as different people may have varying interpretations and decompositions of its functionalities. Therefore, we decided to use file names and counts within the codebase, which are fixed, as one form of ground truth—meaning that if the figure does not contain incorrect files, it can be considered accurate. We compare the files contained in each LLM-generated graph with the file ground truth to approximate the overall accuracy of the map [^2] Accuracy is calculated as $\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ (Baratloo et al., 2015), where TP represents the number of correctly retrieved files. In our context, TN is not applicable since there are no negative files need to be identified. Thus, only TP, FP, and FN are meaningful. Ideally, all files uploaded to the vector store should be analyzed and represented in the graph. The closer the number of files in the graph generated by CodeMap approximates the ground-truth number of files in the codebase, the higher the accuracy should be. .
@@
Results. Results is shown in Figure 7. Overall, additional iterations incrementally improve accuracy. Accuracy stabilizes by the fourth iteration for Python and by the third for Solidity, whereas it has not yet stabilized for the tested Java projects. Moreover, this stabilization trend seems to correlate with codebase size, particularly the number of files: smaller codebases—especially Solidity projects—reach high accuracy after fewer iterations, while larger ones (e.g., Python-668, Solidity-73 and Java-100) require more rounds and exhibit slower gains.
Figure 5. CodeMap Implementation
Best Performance Testing. Since CodeMap ’s prompts still struggle with larger projects, we wonder whether forthcoming models with longer context windows might alleviate this problem. We then simulated such models by running CodeMap with GPT-4.1, which offers a 1 million-token context window (OpenAI, 2025a). We implemented the retrieval pipeline in LangChain because it supports advanced file chunking and filtering [^3] We used OpenAI’s native vector store for file search in previous evaluation because it requires fewer parameters and ensures experimental reproducibility.. Our results showed that through five iterative refinement rounds, CodeMap’s map generation can achieve 99% accuracy on Python-130 and 27% on Python-668. These findings imply that larger context windows can substantially boost CodeMap ’s accuracy; thus, current limitations may be alleviated as LLM’s capability continues to grow.
Figure 6. Prompt structure of business global understanding.
@@ 7. Stage 4: Prototype Validation
To validate the feasibility of CodeMap in real developer settings, we conducted a small-scale observational study with 9 experienced developers and 6 novice developers. The goal was to observe how CodeMap influenced their comprehension process and tool feature usage preference, rather than formal performance comparison. Each participant received a reward equivalent to USD 10–15 for completion of the study.
@@ 7.1.1. Participants
Experienced developer participants (EP). We recruited experienced developer participants from our university and public social media platforms. They included PhD and Master’s students specializing in general software engineering (EP1, EP2), large language models for software engineering (EP3), machine learning (EP4, EP5), and bug mining and vulnerability repair (EP6–EP9). Participants reported 4–10 years of practical programming experience in industry or research projects: EP7–EP9 (6 years), EP6 (7 years), EP4 (8 years), EP2 (4 years), and EP1, EP3, and EP5 (10 years). All participants used ChatGPT daily and primarily programmed in Python, C, C++, and Java. Novice developer participants (NP). For novice developers, we recruited 6 participants (four undergraduate students and two master’s students) from our university’s internal platform to join an in-person study. Their majors included project management (NP1), quantitative finance (NP2), engineering (NP3), data science (NP4), and computer science (NP5, NP6). Participants reported one to five years of programming learning experience (without professional development experience). Most were familiar with Python (NP1–NP4), while some also used Java (NP5) or JavaScript (NP5, NP6). Their ChatGPT usage ranged from monthly (NP3) to weekly (NP1, NP2, NP4) and daily (NP5, NP6).
@@ 7.1.2. Procedure
Experienced participants: Each experienced participant performed short codebase understanding tasks using VSCode (each lasts for 25 minutes), and with the assistance of CodeMap and two baselines: a conversational LLM (i.e., ChatGPT) and a commercial static visualization tool (Understand). When using each tool for codebase understanding, participants were asked to think aloud and to complete short comprehension quizzes and NASA-TLX subjective load questionnaire as reflective probes to prompt purposeful interaction and articulation of understanding and usage experience. Each session lasted approximately 90 minutes, and all interactions were screen recorded for later analysis. Novice participants: For novices, given their limited programming expertise, we did not require them to complete the quiz or use the static analysis tool. Instead, they were asked to explore CodeMap and ChatGPT freely in 10 mins using an example project and to provide feedback.
@@ 7.1.3. Data analysis
To analyse the user’ behaviors, we primarily rely on participants’ experiences through their qualitative verbal feedback, post-study preference questionnaire. For video recordings, we conducted a thematic analysis (Maguire and Delahunt, 2017) of the transcribed data. Two authors collaboratively performed the analysis, engaging in several rounds of team discussions to reach a consensus on interpretation. Moreover, to understand how the system influences users’ behavior, we manually analyzed users’ feature usage time and how their behavior changed within the 25-minute fixed time window for each task. We manually annotated the time spent on each activity through video analysis. In total, we spent 25 hours on video annotation, analyzing 18.48 hours videos.
@@ 7.2. Results
7.2.1. RQ3: How does our expert-inspired system affect experienced and novice developers’ perceived usefulness, intuitiveness, ease of use, and others?
Figure 8. Results: (a) experienced developers’ usability ratings (due to time constraints within the 90-minute task to avoid fatigue, they only rated usability); (b) feature usage time; (c) subjective cognitive load; (d) novice users’ subjective preference.
Experienced Developers. As shown in Figure 8a, all experienced developers (9/9) perceived CodeMap as more intuitive than ChatGPT (2/9) and Understand (3/9). Ease of use ratings were similar between CodeMap (9/9) and ChatGPT (8/9), and considerably higher than Understand (0/9). All participants (9/9) agreed that CodeMap was useful for code understanding, compared with 7/9 for ChatGPT and only 4/9 for Understand.
Novice Developers. Similarly, novice participants rated CodeMap higher across all questions. 4/6 found the tool intuitive, and 5/6 reported it was easy to use. Most novices noted that CodeMap helped them break down information, simplify the code comprehension process, useful in helping understand, etc. However, one participant (EP2), who had a quantitative finance background and had never been exposed to full codebase reading, found CodeMap overwhelming due to the amount of displayed information. She preferred ChatGPT, describing its explanations as clearer and its usability as better than our query feature. This suggests that while CodeMap is promising, it may present a learning curve for users with limited programming experience.
@@ 7.2.2. RQ4: How are experienced and novice developers’ system usage behaviors affected, and why?
Key Finding 1: A shift from text-based to map-based usage occurs when using CodeMap but does not occur when using static visualization tool. We performed a Shapiro–Wilk test on the collected time data and confirmed that all measures followed a normal distribution ($p>.05$). As shown in Figure 8, participants using CodeMap spent significantly less time reading LLM responses (1.15 min) compared with ChatGPT (5.48 min; $t=1.47$, $p=.008$), representing a 79% reduction. Instead, they relied more on viewing system maps to build understanding.
Compared with the visualization baseline (Understand), participants using CodeMap spent 74% less time reading code in VSCode (2.08 min vs. 8.18 min; t=7.87t=-7.87, $p=.0002$) and 90% more time viewing system-provided maps (8.55 min vs. 4.50 min; $t=3.67$, $p=.011$), suggesting that CodeMap ’s natural language–driven visual representations effectively encouraged users to rely more on structured visual information.
@@
Key Finding 2: The shift from text-based to map-based usage occurs because 1) CodeMap can decompose and 2) use natural language to explain modules and their relationships. The high-level map with modules allows users to view the codebase at the module level via semantic descriptions. EP7 described, ”It allows me to intuitively understand the entire project by dividing into modules and shows what each module does. It provides very semantic descriptions, which make it very clear.”
@@
The decomposition not only helps build local understanding but also helps users define a specific context—a challenge to define and describe when using ChatGPT. As EP8 mentioned, ”We usually just upload the entire repo to GPT and let it analyze everything…it definitely loses focus. If we could instead [use CodeMap to] limit its context by selecting specific modules, I think that’s really great.”
@@
In contrast, the information in traditional static analysis tool was not presented appropriately—it lacked key points and textual explanations (EP4, EP5, EP6). The amount of information was overwhelming and difficult to process (EP9), and the call relationships lacked appropriate granularity, making them too complex to understand (EP3). Understand could be suitable for development scenarios where users already built a high-level understanding of the codebase and are familiar with files and module relationships (EP8).
@@
Key Finding 3: The shift from text-based to map-based usage may occur because CodeMap requires less expertise and effort for prompting and may offer greater usefulness in the long term. ChatGPT may only gradually build understanding with already established expertise (EP3). This includes the ability to ask appropriate questions based on the codebase (EP2, EP3, EP7), knowing what they need to know (EP3, EP7, EP8), and using precise questions, including terms from the codebase, to compose their prompts with prompt engineering skills. Like P3 mentioned, ”ChatGPT just analyzes for me whatever I tell it to analyze. So I need to know what I want to analyze.” Without expertise and a clear purpose, the textual explanation can sometimes just list categories without deep interpretation (EP2) or generate answers based on surface-level information, such as file names, rather than deeply understanding the entire content (EP6).
@@
Key Finding 4: CodeMap may demonstrate better effectiveness when applied to large-scale and more complex projects (EP2, EP3, EP4, EP6, EP7). Although our validation study involved relatively small tasks, participants noted that CodeMap would likely be more useful for larger projects. For example, projects with over 10,000 lines of code where ChatGPT often fails to illustrate the relationships between modules (EP7). As EP4 mentioned, ”The reason you’d create diagrams is when a project is really big — when it’s too hard to describe clearly using only language.” Moreover, when projects become larger and more complex, it is difficult for humans to draw high-level maps and enrich with details. As P3 explained, ”If there are a lot of modules, I might roughly draw a diagram [with a few blocks] myself, but I wouldn’t go as deep as your second layer [local understanding].”
@@
Key Finding 5: CodeMap’s understanding flow can decrease novices’ non-purposeful reading and help them build understanding progressively. For example, NP4, an undergraduate student with some basic coding knowledge, felt lost when interacting with ChatGPT and asked questions like ”Can you explain starlette?” [^4] Starlette is a lightweight ASGI (Asynchronous Server Gateway Interface) framework for building high-performance web applications and APIs in Python.. In this lost situation, CodeMap can provide participants a ”starting point”. Like NP5 said, ”This [project architecture understanding guide] is so amazing because when you handle a project for the first time and have no idea what it is about, this is a very good guide to help you understand it.”
@@
Key Finding 6: CodeMap can help novices avoid superficial results from interacting with ChatGPT. Aligning with our formal user study, many novices, when interacting with ChatGPT for analysis, tend to ask simple, straightforward questions, such as the meanings of specific terms (NP4), without understanding it from a big picture. NP6 also mentioned that ChatGPT requires users to actively output and know what to ask. Like he mentioned, ”It will ask, ’What do you want?’ But in CodeMap, a set of standard questions is already present, such as ’What is the business problem about?’, ’What are the relationships between components?’, and ’How does the flow work?’ Everything is already directly presented in CodeMap. If we want to ask more questions, there is the GPT button — you can type and ask about it.” Moreover, with long descriptions from ChatGPT, participants mentioned that the responses were overly descriptive, and they often did not know which parts were correct (NP6).

### v3

@@ 6. Stage 3: Prototype Design
To operationalize the four design opportunities, we designed a prototype, CodeMap.
@@ 6.1. Usage Scenario
We illustrate CodeMap with a concrete scenario in Figure 3, with a demonstration video available on our website. Suppose a novice developer, Bob, needs to understand a new codebase. While reading the source code in VSCode, he uploads the repository to CodeMap. After a brief waiting period, CodeMap generates a high-level business component map of the project. Next, Bob reads the global understanding project overview and follows step-by-step guidance to locate the codebase’s entry point and main modules (Figure 4). Meanwhile, he switches back to VSCode to inspect the original code and verify the relationships shown in the map.
Figure 3. Interface for an example business component map. To support codebase reading, CodeMap allows users to start with 1) Global Understanding through the Global Business Component Map, which provides natural language–enhanced guidance on key modules, components, files, functions, and their relationships. Then, users can 2) zoom into a specific module (e.g., “Security Module”) to view 3) textual instructions in the right pane that describe the module’s details and its relationships with other modules.
@@
Figure 4. An illustration of how users interact with CodeMap to understand a codebase
@@ 6.2. Key Features
  • Global map: business component map: extracting high-level structural information, including business modules, components, and links to show their relationships; function call map: visualizing classes as nodes and showing their inheritance relationships.
@@
  • Natural language enhancement for readability: presenting the links between nodes in the map and natural language to describe the relationships between modules. For example, “Component A provides or manages data for Component B”. Moreover, providing step-by-step guidance for global map to orient users through project’s modules.
In this way, users no longer need to gather information manually and can focus their efforts on understanding the overall structure of the codebase.
Key Feature 2: Seamless switching between layers through interaction.
@@ 6.3. Implementation
Figure 5. CodeMap implementation overview
Figure 6. Prompt structure for global business understanding
@@ 6.4. Accuracy Reliability Checking
Iterative Refinement. We implemented an iterative refinement process–commonly used for LLM-generated outputs (Gu et al., 2024; Liu et al., 2024)—for graph generation; each iteration builds upon the previous iteration’s output. The iteration prompts are available on our website.
Accuracy Evaluation. There is no definitive ground truth for the business component map, as different people may have varying interpretations and decompositions of its functionalities. Therefore, we decided to use file counts within the codebase, which are fixed, as one form of ground truth–meaning that if the generated map contains no incorrect files and fewer missing files, it can be considered more accurate. We compare the files contained in each LLM-generated graph with the file ground truth to approximate the overall accuracy of the map [^4] Accuracy is calculated as $\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ (Baratloo et al., 2015), where TP represents the number of correctly retrieved files. In our context, only TP, FP, and FN are meaningful. TN is not applicable since no files in the codebase are expected to be excluded from the map. Ideally, all files uploaded to the vector store will be analyzed and represented in the graph. The closer the number of files in the graph generated by CodeMap approximates the ground-truth number of files in the codebase, the higher the accuracy should be. .
@@
Results. Results are shown in Figure 7. Overall, additional iterations incrementally improve accuracy. Accuracy stabilizes by the fourth iteration for Python and by the third for Solidity, whereas it has not yet stabilized for the tested Java projects. Moreover, this stabilization trend seems to correlate with codebase size, particularly the number of files: smaller codebases–especially Solidity projects–reach high accuracy after fewer iterations, while larger ones (e.g., Python-668, Solidity-73 and Java-100) require more rounds and exhibit slower gains.
Best Performance Testing. Since CodeMap’s prompts still struggle with larger projects, we wonder whether forthcoming models with longer context windows might alleviate this problem. We then simulated such models by running CodeMap with GPT-4.1555One of the most capable models available when we conducted this experiment in 2025, which offers a 1 million-token context window (OpenAI, 2025a). We implemented the retrieval pipeline in LangChain because it supports advanced file chunking and filtering [^6] We used OpenAI’s native vector store for file search in previous evaluation because it requires fewer parameters and ensures experimental reproducibility.. Our results showed that through five iterative refinement rounds, CodeMap’s map generation can achieve 99% accuracy on Python-130 and 27% on Python-668. These findings imply that larger context windows can substantially boost CodeMap’s accuracy; thus, current limitations may be alleviated as LLM’s capability continues to grow.
@@ 7. Stage 4: Prototype Validation
To validate the feasibility of CodeMap in real developer settings, we conducted a small-scale observational study with 9 experienced developers and 6 novice developers. The goal was to observe how CodeMap influenced their comprehension process and tool feature usage preferences, rather than to conduct a formal performance comparison. Each participant received a reward equivalent to USD 10-15 upon completion of the study.
@@ 7.1.1. Participants
Experienced developer participants (EP). We recruited experienced developer participants from our university and public social media platforms. They included PhD and Master’s students specializing in general software engineering (EP1, EP2), large language models for software engineering (EP3), and machine learning (EP4, EP5), as well as professionals in bug mining and vulnerability repair (EP6-EP9). Participants reported 4–10 years of practical programming experience in industry or research projects: EP7-EP9 (6 years), EP6 (7 years), EP4 (8 years), EP2 (4 years), and EP1, EP3, and EP5 (10 years). All participants used ChatGPT daily and primarily programmed in Python, C, C++, and Java. Novice developer participants (NP). For novice developers, we recruited 6 participants (4 undergraduate students and 2 master’s students) from our university’s internal platform to join an in-person study. Their majors included project management (NP1), quantitative finance (NP2), engineering (NP3), data science (NP4), and computer science (NP5, NP6). Participants reported one to five years of programming learning experience (without professional development experience). Most were familiar with Python (NP1-NP4), while some also used Java (NP5) or JavaScript (NP5, NP6). Their ChatGPT usage ranged from monthly (NP3) to weekly (NP1, NP2, NP4) and daily (NP5, NP6).
@@ 7.1.2. Procedure
Experienced participants: Each participant performed short codebase understanding tasks using VSCode (each lasting 25 minutes), and with the assistance of CodeMap and two baselines: a conversational LLM (i.e., ChatGPT) and a commercial static visualization tool (Understand). When using each tool for codebase understanding, participants were asked to think aloud and to complete short comprehension quizzes and NASA-TLX subjective load questionnaire as reflective probes to prompt purposeful interaction and articulation of understanding and usage experience. Each session lasted approximately 90 minutes, and all interactions were screen recorded for later analysis. Novice participants: For novices, given their limited programming expertise, we did not require them to complete the quiz or use the static analysis tool. Instead, they were asked to explore CodeMap and ChatGPT freely in 10 minutes using an example project and to provide feedback.
@@ 7.1.3. Data analysis
To analyse the user’ behaviors, we primarily rely on participants’ experiences through their qualitative verbal feedback, post-study preference questionnaire. For video recordings, we conducted a thematic analysis (Maguire and Delahunt, 2017) of the transcribed data. Two authors collaboratively performed the analysis, engaging in several rounds of team discussions to reach a consensus on interpretation. Moreover, to understand how the system influences expert users’ behavior, we manually analyzed expert users’ feature usage time and how their behavior changed within the 25-minute fixed time window for each task. We manually annotated the time spent on each activity through video analysis. In total, we spent 25 hours on video annotation, analyzing 18.48 hours videos.
@@ 7.2. Results
7.2.1. RQ3: How does our expert-inspired system affect experienced and novice developers’ perceived usefulness, ease of use, and others?
Figure 8. Results: (a) experienced developers’ usability ratings (due to time constraints within the 90-minute task to avoid fatigue, they only rated usability); (b) and (c) feature usage time; (d) novice users’ subjective preference.
Experienced Developers. As shown in Figure 8a, all experienced participants (9/9) agreed that CodeMap was useful for code understanding, compared with 7/9 for ChatGPT and only 4/9 for Understand. All of them (9/9) perceived CodeMap as more intuitive than ChatGPT (2/9) and Understand (3/9). Ease of use ratings were similar between CodeMap (9/9) and ChatGPT (8/9), and considerably higher than Understand (0/9).
Novice Developers. Similarly, novice participants rated CodeMap higher across all questions. 4/6 found the tool intuitive, and 5/6 reported it was easy to use. Most novices noted that CodeMap helped them break down information, simplify the code comprehension process, useful in helping understand the codebase, etc. However, one participant (NP2), who was from quantitative finance background and had never been exposed to full codebase reading, found CodeMap overwhelming due to the amount of displayed information. She preferred ChatGPT, describing its explanations as clearer and its usability as better than our query feature. This suggests that while CodeMap is promising, it may present a learning curve for users with limited programming experience.
@@ 7.2.2. RQ4: How are experienced and novice developers’ system usage behaviors affected, and why?
Key Finding 1: A shift from text-based to map-based usage occurs when using CodeMap but does not occur when using the static visualization tool. We performed a Shapiro–Wilk test on the collected time data and confirmed that all measures followed a normal distribution ($p>.05$).
As shown in Figure 8b, compared with the visualization baseline (Understand), participants using CodeMap spent 75% less time reading code in VSCode (2.08 min vs. 8.18 min; t=7.87t=-7.87, $p=.0002$) and 90% more time viewing system-provided maps (8.55 min vs. 4.50 min; $t=3.67$, $p=.011$), suggesting that CodeMap’s natural language-driven visual representations effectively encouraged users to rely more on structured visual information.
As shown in Figure 8c, compared with the conversational LLM baseline (ChatGPT) (5.48 min; $t=1.47$, $p=.008$), participants using CodeMap spent significantly less time reading LLM responses (1.15 min), representing a 79% reduction. Instead, they relied more on viewing system maps to build understanding.
@@
Key Finding 2: The shift from text-based to map-based usage occurs because CodeMap can 1) decompose and 2) use natural language to explain modules and their relationships. The high-level map with modules allows users to view the codebase at the module level via semantic descriptions. EP7 described, ”It allows me to intuitively understand the entire project by dividing into modules and shows what each module does. It provides very semantic descriptions, which make it very clear.”
@@
The decomposition not only helps build local understanding but also helps users define a specific context–a challenge to define and describe when using ChatGPT. As EP8 mentioned, ”We usually just upload the entire repo to GPT and let it analyze everything…it definitely loses focus. If we could instead [use CodeMap to] limit its context by selecting specific modules, I think that’s really great.”
@@
In contrast, the information in traditional static analysis tool was not presented appropriately–it lacked key points and textual explanations (EP4, EP5, EP6). The amount of information was overwhelming and difficult to process (EP9), and the call relationships lacked appropriate granularity, making them too complex to understand (EP3). Understand could be suitable for development scenarios where users already built a high-level understanding of the codebase and are familiar with files and module relationships (EP8).
@@
Key Finding 3: The shift from text-based to map-based usage may occur because CodeMap requires less expertise and effort for prompting and may offer greater usefulness in the long term. ChatGPT may only gradually build understanding with already established expertise (EP3). This includes the ability to ask appropriate questions based on the codebase (EP2, EP3, EP7), knowing what they need to know (EP3, EP7, EP8), and using precise questions, including terms from the codebase, to compose their prompts with prompt engineering skills. Like EP3 mentioned, ”ChatGPT just analyzes for me whatever I tell it to analyze. So I need to know what I want to analyze.” Without expertise and a clear purpose, the textual explanation can sometimes just list categories without deep interpretation (EP2) or generate answers based on surface-level information, such as file names, rather than deeply understanding the entire content (EP6).
@@
Key Finding 4: CodeMap may demonstrate better effectiveness when applied to large-scale and more complex projects (EP2, EP3, EP4, EP6, EP7). Although our validation study involved relatively small tasks, participants noted that CodeMap would likely be more useful for larger projects. For example, projects with over 10,000 lines of code where ChatGPT often fails to illustrate the relationships between modules (EP7). As EP4 mentioned, ”The reason you’d create diagrams is when a project is really big — when it’s too hard to describe clearly using only language.” Moreover, when projects become larger and more complex, it is difficult for humans to draw high-level maps and enrich with details. As EP3 explained, ”If there are a lot of modules, I might roughly draw a diagram [with a few blocks] myself, but I wouldn’t go as deep as your second layer [local understanding].”
@@
Key Finding 5: CodeMap’s understanding flow can decrease novices’ non-purposeful reading and help them build understanding progressively. For example, NP4, an undergraduate student with some basic coding knowledge, felt lost when interacting with ChatGPT and asked questions like ”Can you explain starlette?” [^7] Starlette is a lightweight ASGI (Asynchronous Server Gateway Interface) framework for building high-performance web applications and APIs in Python.. Instead, in this lost situation, CodeMap can provide participants a ”starting point”. Like NP5 said, ”This [project architecture understanding guide] is so amazing because when you handle a project for the first time and have no idea what it is about, this is a very good guide to help you understand it.”
@@
Key Finding 6: CodeMap can help novices avoid superficial results from interacting with ChatGPT. Aligning with our formal user study, many novices, when interacting with ChatGPT for analysis, tend to ask simple, straightforward questions, such as the meanings of specific terms (NP4), without understanding it from a big picture. NP6 also mentioned that ChatGPT requires users to actively output and know what to ask. Like he mentioned, ”It will ask, ‘What do you want?’ But in CodeMap, a set of standard questions is already present, such as ‘What is the business problem about?’, ‘What are the relationships between components?’, and ‘How does the flow work?’ Everything is already directly presented in CodeMap. If we want to ask more questions, there is the GPT button — you can type and ask about it.” Moreover, with long descriptions from ChatGPT, participants mentioned that the responses were overly descriptive, and they often did not know which parts were correct (NP6).